OpenStudy (anonymous):

R= MILE RUN & B=MILE BIKED 7R+4B=90 8R+3B=84 HOW MANY MILE WILL JOHN RUN DURING THE RACE?

OpenStudy (anonymous):

I WAS TRYING TO DO IT AND I GOT R=6.0 AM I RIGHT

OpenStudy (anonymous):

6.00000001=R

OpenStudy (anonymous):



OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

jello

OpenStudy (anonymous):

Just checked over everything and it looks good. Btw, if you keep -4*8/3 as -32/3, you will find that it is exactly 6. So @ 7R + 4( 28 - (8/3)R) =90. It goes (21/3)R + 112 - (32/3) R = 90, -(11/3)R = -22, -11R = -66, and finally R = 6.

OpenStudy (anonymous):

YEEEEEESS THANK U